OpenStudy (anonymous):

what is the slope of the line passing through points A(-2,5) and B(9,-4)

OpenStudy (anonymous):

The formula to find the slope between two points is m = y2 - y1/x2 - x1

OpenStudy (anonymous):

My teacher gave me the answer it's -9/11 I just don't know how to show my work

OpenStudy (anonymous):

m = (-4) - 5/9 - (-2) m = -9/11

OpenStudy (anonymous):

thnx so much

OpenStudy (anonymous):

no problem(: